Design a PLA like a ROM. A programmable logic array (PLA) performs the same
Answer #1

Inputs A В AB +V +V +V AC BC AC +V 0 F1 F Outputs Fig 9 -PL A with 3 inputs, 5 product t ems and 4 outputsFO 2n(0,1,4,6)- AB+AC F1 Σ1n(2.3.4.6.7) = B+AC F2 - Zm(0,1,2,6) -AB+BC F3- Žm(2,3,5,6,7) AC+B im m(2,5 The above set ofOR AND Fig 10-AND-OR array equivalent of Fig. 9 The contents of a PLA can be specified by a modified truth table as shown inInputs Outputs ProductA B CIFO F1 F2 F3 AB AC 0 1 0 1 BC AC fig II - PLA Table for Fig. lt0 The first row of the table indicF1=Σm(2,3,5,7,8,9, 10,11,13,15) F2= Σm(2,3,5,6,7,10,11,14,15) F3-m(6,7,8,9,13,14,15) (1) abd abc ab ab ab cd 00 01 1 1 Λ0/ cPLA Realization of Equations Inputs abd ab bc Word Lines bc Outputs Fig 14-PLA Realization of Equations 7, 8 and 9. Fig 14